kind of confused on this:
suppose the functions f and g are defined for all real numbers x as follows: f(x)=3x^2
g(x)=2x
write the expressions for (f+g)(x)and (f-g)(x) and evaluate (fxg)(-2)

Answer by stanbon(75887)

suppose the functions f and g are defined for all real numbers x as follows: f(x)=3x^2
g(x)=2x
write the expressions for
(f+g)(x) = 3x^2 + 2x
-----------------------------------
(f-g)(x) = 3x^2-2x
---------------------------
(fxg)(-2) = f(-2)*g(-2) = 3(4) + 2(-2) = 12-4 = 8
------
Cheers,
Stan H.
-------------